OpenStudy (anonymous):

solve the following systems of equations show work please. -x+8y=16 3x+4y=36

jimthompson5910 (jim_thompson5910):

-x+8y=16 3x+4y=36 3(-x+8y)=3*16 3x+4y=36 -3x+24y=48 3x+4y=36 ---------- 0x+28y=84 28y=84 y=84/28 y=3 -x+8y=16 -x+8(3)=16 -x+24=16 -x=16-24 -x=-8 x=8 So the solutions are x=8 and y=3 giving us the ordered pair (8,3)

hero (hero):

y = (16+x)/8 y = (36-3x)/4 y = y 4(16+x) = 8(36-3x) 64+4x = 288 - 24x 28x = 224 x = 8 y = 3